# Toward Routing River Water in Land Surface Models with Recurrent Neural Networks

### v1

@@ 2.1 From precipitation in the USA to runoff worldwide
This series of experiments investigates the performance of LSTMs trained and validated using basins in the USA only and LSTMs trained and validated on global data. The USA was chosen as reference for being a well characterized region in terms of data availability and because it allows for a more direct comparison to previous results (Kratzert et al., 2019a, b). Our goals were to determine how the performance of the model changes when the dynamic input changes (from precipitation to runoff driven from reanalysis), to determine how the model performance changes when trained with more varied data (from the USA to the global case), and to investigate how the model performs at the basin and temporal generalization tasks.
2.1.1 Temporal generalization
To address these questions, we begin with a temporal training/validation split. The dataset was divided into three different timeseries for all basins: from 01/10/1999 to 30/09/2009 for training, from 01/10/2009 to 30/09/2019 for validation, and from 01/10/1989 to 30/09/1999 for testing. We compared the model trained on the USA with runoff input with the model trained on the USA with precipitation input. We also compared the performance of the model trained on the USA with runoff input with the model trained using the global data set with runoff input.
@@ 2.1.3 Summary of results
2.2 Comparison with physics-based model
In a second series of experiments, we compare the LSTM model with the LISFLOOD streamflow model (Van Der Knijff et al., 2010), provided by GloFAS-ERA5 (Harrigan et al., 2020). This choice was made mainly because, like LISFLOOD, our streamflow results are based on training a routing model with runoff from ERA5-Land as input. Moreover, the comparison between the LSTM and LISFLOOD was easy to do without additional calibrations, as the LISFLOOD streamflow predictions are provided with the reanalysis data.
That said, there are some important differences between our setting and that of GloFAS-ERA5: (i) LISFLOOD is a complete river routing model, whereas the LSTM models basins independently; (ii) the objective function for LISFLOOD is based on the KGE metric whereas the LSTM was trained to optimize NSE; and (iii) LISFLOOD is calibrated with a different global set of gauges. However, most of those gauges come from GRDC, as stated in Alfieri et al. (2020). When considering the validation time frame of the LISFLOOD model, it may coincide with the test period outlined in this paper, potentially simplifying the prediction task. However, we cannot be sure of this overlap. Hence, we presume the model to be configured more akin to a time-split configuration. To facilitate the comparison, we made an effort to provide a one-to-one relation between the LSTM/time-split training set and GloFAS-ERA5 by doing the following:
We consider only gauged basins that are independent of other basins, restricting the analysis to the set of basins used by our global model in the all-basins, time-split configuration;
We filter basins by allowing no more than a 20% difference between the catchment area of GRDC and the upstream area for the associated point from GloFAS, which we chose to be the grid point containing the gauge;
We filter basins by selecting only basins for which the corresponding grid cell in GloFAS lies entirely inside the basin polygon limits.
The process resulted in a total of 639 basins in the time-split and 190 basins in the basin-split configurations, which is about half the basins used in the global dataset. Using these basins, we evaluate the LSTMs in both time-split and basin-split configurations, after training with the previously described global data with no additional calibration. The LSTMs are evaluated against GloFAS-ERA5 for the corresponding time periods.
@@ 2.2 Comparison with physics-based model
For the gauges we use, the LSTM in the time-split configuration shows an overall better accuracy compared with GloFAS, while the LSTM in the basin-split configuration displays a worse median accuracy in both metrics. For the NSE, the LSTM has a median score of 0.61 in time-split and 0.34 in basin-split configuration, whereas GloFAS displays a median score of 0.39. For the KGE, the LSTM has a median score of 0.71 in time-split and 0.43 in basin-split configuration, whereas GloFAS displays a median score of 0.54. In both cases, GloFAS (which, as we highlighted, is believed to be in a configuration closer to a time-split) is quantitatively closer to the LSTM basin-split configuration. This highlights the capabilities of the LSTM model, since the basin-split is a more difficult configuration. We show other metrics in Table 3.
@@
LSTM basin-split | 25.26% | 0.49 | 0.34 | 11.05% | 0.42 | 0.43 | |
LSTM time-split | 7.98% | 0.59 | 0.61 | 2.66% | 0.62 | 0.71 | |
GloFAS-ERA5 | 24.1% | 0.50 | 0.39 | 11.27% | 0.52 | 0.54 |
The LSTM models perform better here compared with when evaluated on the data in section 2.1. This may suggest that the set of basins used to compare with the GloFAS benchmark contains gauges with more reliable measurements.
@@
In this section, we present some simulated timeseries along with observations for various values of NSE and KGE to get a visual sense of performance. For four different basins, Figure 6 shows the predictions of the LSTM model in the global time-split configuration, the GloFAS-ERA5 reanalysis, and the observed streamflow at the corresponding GRDC gauge.
@@ 2.4 Geographic patterns
We also investigate the aridity index, which is provided in the HydroATLAS dataset (Linke et al., 2019) and has been found to be a strong predictor of the model’s performance. Figure 8 shows a trend toward poorer performance in drier regions (i.e., regions with lower aridity index), consistent with findings in Feng et al. (2020), but now on a global scale. (For context, the basin depicted in Figure 6a has an aridity index of 0.26.) This trend extends to other variables, such as mean runoff, for which the model also shows poorer performance in drier regions.
@@ 3 Discussion and conclusion
Long short-term memory models have been shown to be the state of the art for modeling streamflow in hydrological systems (Kratzert et al., 2019b), but most studies using these models have so far been restricted to specific regions, and they have focused on representing the entire land hydrological system as a single entity (Kratzert et al., 2019b; Koch and Schneider, 2022). However, for integration into the land surface models used in climate modeling, it is crucial for river models to demonstrate proficiency in generalizing across basins and temporal scales globally, using both surface and subsurface runoff data, rather than, for example, precipitation data. This necessity arises because traditional river models in land surface models merely route water, leaving the modeling of snow, soil, and vegetation hydrological processes to other model components. In this study, we have taken concrete steps toward developing an ML model that can substitute traditional river routing models within LSMs. We have successfully trained and validated an LSTM at the task of predicting streamflow from runoff worldwide.
We began our analysis by contrasting the results from an observed precipitation-driven model, which holistically represents land hydrology, with those of a modeled runoff-driven model. For the United States, we found that an LSTM trained only to route runoff performs comparably to one designed to simulate the entire land hydrology system (Kratzert et al., 2019b), even when trained on potentially less accurate and biased runoff data. A possible drawback of our approach is that it relies on modeled runoff from reanalysis as training data, which likely is less accurate than using precipitation data directly. Recently, other studies have used precipitation data on a global scale to simulate streamflow from LSTMs, being able to outperform the GloFAS model even when trained in a basin-split configuration, similar to the one made on this work (Nearing et al., 2024). However, it is not clear if this increase in performance is due to a change in the LSTM model, training data, or due to the forcing data used.
@@
Our exploration into the model’s generalization capabilities revealed that a globally trained model, as opposed to one trained exclusively using data from the USA, achieved superior performance when both were fed runoff data in a time-split configuration. This improvement underscores the potential benefits of incorporating diverse global data. However, when evaluating models trained in the basin-split configuration, the performance gain was not as pronounced, highlighting the complex challenge of global basin generalization. This challenge is exacerbated by imbalances in global data availability and the wide range of possible climate conditions across the globe, which may not be well sampled in geographically localized training data (e.g., from the USA).
In comparison to the reanalysis data from the physics-based LISFLOOD model (GloFAS-ERA5), our LSTM model showed more accurate forecasts in a gauged scenario (time-split) but fell short in an ungauged scenario (basin-split). We argued that results from GloFAS-ERA5 are closer to a gauged (time-split configuration) scenario than an ungauged (basin-split) scenario. Thus, these findings suggest that our LSTM model holds promise for global basin generalization, despite the discrepancies between scenarios.
@@
Additionally, our analysis revealed a correlation between the model’s performance and the aridity index, suggesting that drier regions pose unique challenges for the LSTM model. This could be due to the model’s difficulty in capturing short-term precipitation events that trigger streamflow peaks in these areas, as seen in Figure 6a (Feng et al., 2020). It also underperformed in regions likely underrepresented in the training data. Despite these challenges, the model’s time generalizability was consistent across different basin sizes, even if better basin generalizability was observed for bigger basins, as discussed in Appendix C. We highlight that the model doesn’t need to be re-calibrated for the evaluation of different basin levels.
In conclusion, our study presents a promising step toward using neural networks for water routing in the LSM component of climate models. The promising results motivate further research to refine these models for comprehensive river routing and integration into climate models. Future efforts will focus on designing training strategies to mitigate error propagation across connected basins and incorporating architectural adjustments to ensure mass conservation (e.g., (Hoedt et al., 2021)), potentially enhancing long-term simulation accuracy.
@@
All data used to generate our training and test data are open source. The source code for the data engineering, the models, the extraction from the benchmark, and the visualizations can be found in https://github.com/maumlima/Rivers.
@@ Appendix A List of static attributes used in the models
Appendix B Hyper-parameters
Table 6 lists the most important hyper-parameters used in the model in the global time-split configuration. The learning rate started at and gradually decreased to during the 35 epochs of training.
Hyper-parameter | Value |
[…]
Droput rate | 0.4 |
Length of input sequence | 270 |
[…]
@@ Appendix D Mass balance
If there is no loss of water due to infiltration into deeper layers in the ground or evaporation into the atmosphere, the sum of runoff over the area of a catchment should match the discharge at the outlet when averaged over extended periods. We calculate the absolute difference between the outflow from each basin (normalized by its area) and the total runoff within the corresponding catchment over a 9-year time window. The results of this calculation are presented in Figure 10 for streamflow (in blue): forecasted by the LSTM (Figure 10a), recorded by observations from GRDC gauges (Figure 10b), and provided by the GloFAS reanalysis (Figure 10c). In each case, the number of outliers (“N of outliers”) represents the quantity of data points falling outside the boundaries of the graphs. These results indicate that the LSTM conserves mass at a level comparable to the observations and a physical model. While this is a positive result, it implies that additional processes like evaporation from rivers, re-infiltration of water into the ground, and human interference may need to be understood and modeled in order to achieve a closed water balance. One could expect to find a better mass conservation from the physics-based model, but it should be noted that GloFAS has parameters which control mass loss (to underground storage, for example) and the actual state of mass conservation can vary depending on the version of runoff data that was used in their calculations.

### v3

@@ 2.1 From precipitation in the USA to runoff worldwide
This series of experiments investigates the performance of LSTMs trained and validated using basins in the USA only and LSTMs trained and validated on global data. The USA was chosen as reference for being a well characterized region in terms of data availability and because it allows for a more direct comparison to previous results (Kratzert et al. (2019a, b); see Section 2.2.2 for further discussion). Our goals were to determine how the performance of the model changes when the dynamic input changes (from precipitation to runoff driven from reanalysis), to determine how the model performance changes when trained with more varied data (from the USA to the global case), and to investigate how the model performs at the basin and time generalization tasks.
2.1.1 Time generalization
To address these questions, we begin with a time training/validation split. The dataset was divided into three different timeseries for all basins: from 01/10/1999 to 30/09/2009 for training, from 01/10/2009 to 30/09/2019 for validation, and from 01/10/1989 to 30/09/1999 for testing. We compared the model trained on the USA with runoff input with the model trained on the USA with precipitation input. We also compared the performance of the model trained on the USA with runoff input with the model trained using the global dataset with runoff input.
@@ 2.1.3 Summary of results
2.2 Comparison with other models
2.2.1 Comparison with a physics-based model
In a second series of experiments, we compare the LSTM model with the LISFLOOD model, provided by GloFAS v4.0. This choice was made because LISFLOOD explicitly calculates runoff and performs routing of river water from reanalysis data from ERA5 (Hersbach et al., 2020). Moreover, the comparison between the LSTM and LISFLOOD was straightforward to carry out without additional calibrations, as the LISFLOOD streamflow predictions are provided with the reanalysis data. That said, there are some important differences between our model and training procedure and those of GloFAS: (i) GloFAS uses different forcings from ERA5, including precipitation, to calculate runoff internally, whereas the LSTM uses runoff variables computed by the HTESSEL model (Balsamo et al., 2009) in ERA5-Land (ii) LISFLOOD is a complete river routing model, whereas the LSTM models basins independently; (iii) the objective function for GloFAS is based on the KGE metric whereas the LSTM was trained to optimize the NSE; and (iv) GloFAS is calibrated with a different global set of gauges.
Below, we compare the results of GloFAS to the LSTM in both a time- and basin-split experiment. However, several caveats are necessary to make. First, we do not know the exact time-split used in the GloFAS calibration. This means that the validation time frame of the LISFLOOD model may coincide with the test period used in our comparison here, potentially simplifying the prediction task of the time-split experiment for GloFAS relative to the LSTM. Second, the locations of only GRDC gauges that were used to calibrate GloFAS are known and other datasets may have been used in place of the same gauges (GloFAS team members, private communication). This allows for a basin-split experiment comparison, but (a) some of the gauges in our test set may have been used in the calibration of GloFAS, and (b) the unknown time-split may mean that for some of the test basins, GloFAS is generalizing in both space and time. The former makes the GloFAS prediction task easier than the LSTM task, while the latter makes the GloFAS prediction task harder than the LSTM task, since the LSTM is trained and tested in different basins but using the same time window. For a large enough temporal training period and for a global set of basins, we would expect that generalizing in space and time is similar to generalizing in space, since that is the much more challenging task. For all of these reasons, we note the basin- and time-split experiments with a "*" in the Table and Figure presented in this Section.
In order to construct the set of gauges used in our comparisons, we considered the following:
When carrying out the time-split experiment, we restrict the analysis to the intersection of the test set of basins used by our global time-split model and the basins used to calibrate GloFAS.
When carrying out the basin-split experiment, we restrict the analysis to the intersection of the test set of basins used by our global basin-split model and the basins not used to calibrate GloFAS.
We filter basins by allowing no more than a 20% difference between the catchment area reported by GRDC and the catchment reported by GloFAS.
The process resulted in a total of 283 basins in the time-split and 197 basins in the basin-split configurations. We highlight that these basins are independent of other basins, since our models do not route water between basins. The LSTMs are evaluated against GloFAS on the same time periods.
@@ 2.2.1 Comparison with a physics-based model
For the gauges we use, the LSTM shows an overall better accuracy compared with GloFAS in both basin- and time- split experiments. For the temporal split, the LSTM has a median score of 0.64 on NSE and 0.72 on KGE, whereas GloFAS displays a median score of 0.53 on NSE and 0.71 on KGE. For the basin split, the LSTM has a median score of 0.33 on NSE and 0.41 on KGE, whereas GloFAS displays a median score of 0.30 on NSE and 0.40 on KGE. Observe that despite being optimized for NSE, the LSTM has superior scores on both basin- and time-split to GloFAS on KGE. Moreover, the LSTM predictions result in fewer basins with performance worse than the baseline in all experiments. We show other metrics in Table 3.
@@
| LSTM basin-split | 26.4% | 0.51 | 0.34 | 12.18% | 0.44 | 0.41 | |
| GloFAS basin-split* | 34.01% | 0.45 | 0.30 | 13.2% | 0.41 | 0.40 | |
| LSTM time-split | 7.07% | 0.62 | 0.64 | 1.77% | 0.66 | 0.72 | |
| GloFAS time-split* | 15.19% | 0.58 | 0.56 | 6.71% | 0.66 | 0.71 | |
It is interesting that the LSTM time-split model performs better here compared with that of Section 2.1. This may suggest that the set of basins used by GloFAS contains gauges with more reliable measurements.
2.2.2 Comparison with other LSTM models
In this section we discuss possible comparisons to similar LSTM-based models from Kratzert et al. (2019b) and Nearing et al. (2024). These comparisons are in general difficult to make as each model is trained using different basins, using different dynamic input variables, and, in the case of Nearing et al. (2024), with different experimental designs. The models are also intended for different use cases.
In Kratzert et al. (2019b), an LSTM model was trained using observed dynamic inputs gathered for catchments in the CONUS (Addor et al., 2017). In a time-split experiment, Kratzert et al. (2019b) reported a median NSE of 0.73 (without ensembles). In our time-split experiment for the USA with reanalysis precipitation as input (Section 2.1), we found a median NSE of 0.52. Importantly, the dynamic inputs of the former come from observations, while the dynamic inputs used to train the latter come from reanalysis data. As a consequence, these two models are not easily comparable as using observed precipitation (among other inputs) is more accurate than reanalysis data. Additionally, our LSM-RNN was calibrated with fewer basins.
In Nearing et al. (2024), an LSTM model was trained using a rich set of dynamic inputs coming from multiple datasets including high-resolution forecasts, estimates from satellites, and ERA5-Land reanalysis data. Precipitation, and not runoff, was used. The purpose of the model is for forecasting floods in watersheds globally. The model was training globally, and was tested in a time-and-basin split simultaneously. Due to large differences in input data and experimental design, a comparison between the published results of this model and our model is not necessarily meaningful to make. Using more input data with higher accuracy than reanalysis data should improve streamflow predictions, and hence we expect this model to outperform our LSM-RNN.
@@ 2.2.2 Comparison with other LSTM models
In this section, we present some simulated timeseries along with observations for various values of NSE and KGE to get a visual sense of performance. For four different basins, Figure 6 shows the predictions of the LSTM model in the global time-split configuration, the GloFAS reanalysis, and the observed streamflow at the corresponding GRDC gauge.
@@ 2.4 Geographic patterns
We also investigate the aridity index, which is provided in the HydroATLAS dataset (Linke et al., 2019) following the dataset of Zomer et al. (2008). The aridity index is given by the ratio between the mean annual precipitation and the mean annual evapotranspiration, and is calculated on a per grid cell basis. Hence, lower values of aridity index represent arid climates and higher values represent humid climates, for further discussion, one may reference to the newest version of the dataset in Zomer et al. (2022). In our analysis, the aridity index has been found to be a strong predictor of the model’s performance. Figure 8 shows a trend toward poorer performance in drier regions (i.e., regions with lower aridity index), consistent with findings in Feng et al. (2020), but now on a global scale. (For context, the basin depicted in Figure 6a has an aridity index of 0.26.) This trend extends to other variables, such as mean runoff, for which the model also shows poorer performance in drier regions.
@@ 3 Discussion and conclusion
Long short-term memory models have been shown to be the state of the art for modeling streamflow in hydrological systems (Kratzert et al., 2019b), but most studies using these models have so far been restricted to specific regions, and they have focused on representing the entire land hydrological system as a single entity (Kratzert et al., 2019b; Koch and Schneider, 2022). However, for integration into the land surface models used in climate modeling, it is crucial for river models to demonstrate proficiency in generalizing across basins and time scales globally, using surface and subsurface runoff data, rather than precipitation data. This necessity arises because traditional river models in land surface models route water, but leave the modeling of snow, soil, and vegetation hydrological processes to other model components. In this study, we have taken concrete steps toward developing an ML model that can substitute traditional river routing models within LSMs.
@@
In that effort, we have trained and validated an LSTM at the task of predicting streamflow from modeled runoff worldwide. We began our analysis by contrasting the results from an reanalysis precipitation-driven model, which holistically represents land hydrology, with those of a modeled runoff-driven model. For the United States, we found that an LSTM trained only to route runoff performs comparably to one designed to simulate the entire land hydrology system, even when trained on potentially less accurate and biased runoff data. Our exploration into the model’s generalization capabilities revealed that a globally trained model, as opposed to one trained exclusively using data from the USA, achieved superior performance when both were fed runoff data in a time-split configuration. This improvement underscores the potential benefits of incorporating diverse global data. However, when evaluating models trained in the basin-split configuration, the performance gain was not as pronounced, highlighting the complex challenge of global basin generalization. This challenge is exacerbated by imbalances in global data availability and the wide range of possible climate conditions across the globe, which may not be well sampled in geographically localized training data (e.g., from the USA).
@@
Additionally, our analysis revealed a correlation between the model’s performance and the aridity index. While streamflow in arid basins can be modeled well by the LSTM, it is also true that all basins with a poor NSE have a lower aridity index. This suggests that drier regions pose challenges for the LSTM model, but that other basin features may affect performance as well. The dependence on aridity could be due to the model’s difficulty in capturing short-term precipitation events that trigger streamflow peaks in these areas, as seen in Figure 6a (Feng et al., 2020). It also underperformed in regions likely underrepresented in the training data. Despite these challenges, the model’s time generalizability was consistent across different basin sizes, even if better basin generalizability was observed for bigger basins, as discussed in Appendix C. We highlight that the model doesn’t need to be re-calibrated for the evaluation of different basin levels, which correspond to different spatial resolutions of LSMs.
We also took steps to compare our model performance with the performance of others. In general, this is challenging due to differences in experimental design, loss function choice, and data used. These differences are motivated by differences in intended application of the model, and hence these comparisons are not always meaningful to make. That said, we would like to highlight two main conclusions we drew from our comparisons.
Using a similar dataset, experimental design, and LSTM architecture, we found that a model trained with reanalysis data (our USA time-split precipitation model) had a significantly worse performance compared to the LSTM of Kratzert et al. (2019b) which was trained with observations (a drop in median NSE from 0.73 to 0.52). This indicates that modeled variables (e.g., precipitation or runoff from reanalysis) suffers from larger errors than observed precipitation, which is not surprising, and that these errors affect the quality of the predicted streamflow. An LSTM trained using modeled runoff - for which observations are not possible - may learn both to predict streamflow and to correct biases in modeled runoff. However, for our use case of a river model routing water within an LSM, and not modeling the entire land hydrology system, this may be unavoidable.
In contrast, the physics-based LISFLOOD model (GloFAS) is more similar to our model in that it uses modeled runoff (calculated internally) as a dynamic input to perform streamflow routing. We have compared the models on a subset of the gauges used to calibrate GloFAS (time-split) and on a subset of gauges not used for calibration (basin-split). The LSTM model showed more accurate simulations in both scenarios, but note that there are still some differences in experimental design between the two. Unfortunately, a recalibration of the model using our data and experimental design is beyond the scope of this work – previous works have reported calibration of GloFAS that took months in computational time (Alfieri et al., 2020).
In order to integrate our model in with an LSM, it must be extended to channel routing and it must respect physical priniciples like mass conservation. To our knowledge, no article has demonstrated a machine learning-based approach for global routing simulations, although some approaches show how to account for inter-basin connections in a restricted region (Moshe et al., 2020). On the other hand, physical models such as LISFLOOD model routing between basins and hence are currently the standard for LSMs. One may interpret this work as the hillslope routing component of a complete routing model. In that case, two possible ways to extend this to route water between basins are to use a physics-based channel routing model or to use similar neural networks to additionally perform river channel routing. For the latter, designing training strategies to mitigate error accumulation across connected basins may be necessary. Moreover, incorporating architectural adjustments to ensure mass conservation (e.g., Hoedt et al., 2021), is a necessary step for coupling with climate models, potentially enhancing long-term simulation accuracy. In conclusion, we have demonstrated that machine learning based river models exhibit basin and time generalizability, a requirement for use in a global climate model. Our study presents the first step toward using these models in the LSM component of climate models and our results motivate further research to extend these models for comprehensive river routing.
@@
All data used to generate our training and test data are open source. The source code for the data engineering, the models, the extraction from the benchmark, and the visualizations can be found in https://github.com/limamau/Rivers.
@@ Appendix A List of static attributes used in the models
Appendix B LSTM architecture
As explained in Section 1.2, the LSTM model is a recurrent neural network, where one cell is used recursively for a sequence length of iterations. In the text, we have represented this cell as a flexible parameterized model . The equations that describe the internal mechanism of such a function (i.e., a forward pass in an LSTM cell at the iteration) are:
where , , and are the input, forget, cell and output gates, respectively. Each of these gates have their learnable weights for both the inputs (represented by the matrix and hidden states (represented by the matrix , as well as learnable biases (represented by the vector ). Observe that these learnable parameters don’t depend on time – even if the result of each gate may depend on time with and , they use the same learnable parameters, which are applied recursively. The input, forget and output gates are enclosed by the sigmoid function and the cell gate is enclosed by the hyperbolic tangent function . Furthermore, in the expression of cell state and the hidden state , represents an element-wise multiplication. Figure 9 schematizes these equations inside the LSTM cell.
Observe that, in comparison to Figure 3, we explicitly represent the cell state , which is also a “hidden" state that influence subsequent iterations. The general idea of the LSTM is to add or remove information of this cell state through gates, which is what is written in (10). On the left side of the equation (), the forget gate (which is enclosed by a sigmoid function) can remove (resp. add) information from the cell state by multiplying each element of the vector by 0 (resp. 1). On the right side of the equation (), the input gate applies the same logic to the cell gate to remove or add information to the cell state. For more intuition behind this architecture, we point out to the comprehensive text in https://colah.github.io/posts/2015-08-Understanding-LSTMs/ (last accessed 13 August 2024).
@@ Appendix D Mass balance
If there is no loss of water due to infiltration into deeper layers in the ground or evaporation into the atmosphere, the sum of runoff over the area of a catchment should match the discharge at the outlet when averaged over extended periods. We calculate the absolute difference between the outflow from each basin (normalized by its area) and the total runoff within the corresponding catchment over a 9-year time window. The results of this calculation are presented in Figure 11 for streamflow (in blue): simulated by the LSTM (Figure 11a), recorded by observations from GRDC gauges (Figure 11b), and provided by the GloFAS reanalysis (Figure 11c). In each case, the number of outliers (“N of outliers”) represents the quantity of data points falling outside the boundaries of the graphs. We highlight that the runoff calculated internally by LISFLOOD was not used here, but rather the one provided by ERA5-Land, which was calculated by HTESSEL – a different model forced with similar ERA5 data. Under these circumstances, the LSTM conserves mass at a level comparable to the observations and a physical model. While this is a positive result, it implies that additional processes like evaporation from rivers, re-infiltration of water into the ground, and human interference may need to be understood and modeled in order to achieve a closed water balance. One could expect to find a better mass conservation from the physics-based model, but it should be noted that GloFAS has parameters which control mass loss (to underground storage, for example) and the actual state of mass conservation can vary depending on the version of runoff data that was used in their calculations.
@@
ML, KD, ORAD and TS: conceptualization, formal analysis, visualization, writing.
The authors declare no competing interests.